# Bidirectional Uncertainty-Aware Region Learning for Semi-Supervised Medical Image Segmentation

### v1

@@ Abstract
In semi-supervised medical image segmentation, the poor quality of unlabeled data and the uncertainty in the model’s predictions lead to models that inevitably produce erroneous pseudo-labels. These errors accumulate throughout model training, thereby weakening the model’s performance. We found that these erroneous pseudo-labels are typically concentrated in high-uncertainty regions. Traditional methods improve performance by directly discarding pseudo-labels in these regions, but this can also result in neglecting potentially valuable training data. To alleviate this problem, we propose a bidirectional uncertainty-aware region learning strategy. In training labeled data, we focus on high-uncertainty regions, using precise label information to guide the model’s learning in potentially uncontrollable areas. Meanwhile, in the training of unlabeled data, we concentrate on low-uncertainty regions to reduce the interference of erroneous pseudo-labels on the model. Through this bidirectional learning strategy, the model’s overall performance has significantly improved. Extensive experiments show that our proposed method achieves significant performance improvement on different medical image segmentation tasks.
@@ 1 Introduction
With the continuous development of deep learning, many medical image segmentation tasks [1, 2] have achieved great success through fully supervised learning using large amounts of labeled data. However, it is challenging to obtain large-scale and accurately labeled medical datasets to train segmentation models due to the privacy, dispersion, and labeling difficulties of medical images [3]. Semi-supervised medical image segmentation [4, 5, 6, 7] effectively reduces the cost of labeled data and improves the performance and generalization ability of the model to a certain extent by utilizing a small amount of labeled data and a large amount of easily available unlabeled data to train the model.
Figure 1: We visualize the regions of pseudo-label errors relative to the real label and the regions of high entropy predicted by the model and find that most of the regions of pseudo-label errors overlapped with the regions of high entropy predicted by the model. We also visualize category prediction probabilities using reliable and unreliable predictions. Specifically: (a) Pseudo-label errors relative to real label, (b) Selecting predicted high entropy regions above the 99th percentile, (c) Category-wise probability of a reliable prediction, which is confident enough for supervising the class right ventricle, (d) Category-wise probability of an unreliable prediction, which hovers between right ventricle, Myocardium, and Left Ventricle, yet is confident enough of not belonging to background.
@@
In semi-supervised medical image segmentation, effectively leveraging unlabeled data is crucial [8]. A common approach is assigning pseudo-labels to unlabeled data [9, 10], allowing the model to learn from these approximate “ground-trurh”. However, generating high-quality pseudo-labels becomes particularly difficult due to the scarcity of labeled data. Low-quality pseudo-labels introduce greater uncertainty and contain numerous errors, which can mislead the model during training and ultimately affect its learning outcomes and performance [11]. Therefore, effectively preventing the accumulation and propagation of errors in pseudo-labels has become a critical challenge in this area of research. Some researchers have proposed predictive filtering based on confidence scores to solve this problem. They argue that pseudo-tags generated by predictions with higher confidence scores are more reliable [12]. This approach aims to filter out low-confidence predictions and to retrain the model using only high-confidence pseudo-labels. However, one potential problem caused by only using reliable predictions is that some pixels may never be learned in the entire training process, resulting in inadequate model training [13]. In addition, determining a threshold that effectively distinguishes between high-confidence and low-confidence pseudo-labels is also extremely challenging. Despite the ability of dynamic thresholds [14, 15] to adapt automatically to different disturbance conditions, allowing the model to select pseudo-labels during training flexibly, their effectiveness relies on the model’s current state. Suppose the model’s prediction is conservative at a certain moment in time. In that case, it will set a higher confidence threshold, thus also leading to the exclusion of some valuable low-confidence samples. Since each region in a medical image may contain critical medical information [16], retaining only high-confidence predictions may result in the loss of critical information, which may lead to the model’s inability to globally capture subtle differences in the medical image, reducing its ability to recognize complex structures.
To address this issue, we first analyze the model’s predicted entropy maps and the pseudo-label error maps obtained by comparing pseudo-labels with ground-truth labels, as shown in Fig. 1 (a) and (b), we find that incorrect pseudo-labels often occur in high-entropy regions, which largely coincide with the boundaries of the segmented targets. Therefore, we can effectively detect and correct errors in the pseudo-labels by identifying and processing these high-entropy regions. Additionally, as shown in Fig. 1 (c) and (d), although the model’s predictions on the high-entropy regions are unreliable, it is very confident that these pixels do not belong to the background, effectively utilizing this information can enhance the model’s learning. Based on these observations, we identified that the core issue lies in effectively locating and leveraging uncertain regions. Therefore, we propose a new bidirectional region learning strategy for semi-supervised learning, which includes Uncertainty-aware Region Learning (URL) and Certainty-aware Region Learning (CRL). Unlike previous approaches that remove erroneous pseudo-labeled regions, we identify uncertain regions by calculating the entropy of the model’s predictions and dynamically adjust the learning weights of these regions based on the model’s prediction results to address these challenges. Specifically, for each input image, we first generate predictions through the model and then generate the corresponding entropy value matrix based on the model predictions. The entropy value matrix evaluates the certainty of the predicted pixels in different categories. Then, we classify the predictions into high-entropy and low-entropy regions, i.e., uncertain and certain regions, based on the entropy value matrix. To motivate the model to focus on learning these potentially uncertain regions, we assign higher learning weights to the uncertain regions during the supervised learning phase of the labeled data, since accurate labeling allows the model to maximize the learning and correct the segmentation errors in these regions. Meanwhile, to reduce the negative impact of uncertain regions on model performance, in the self-supervised learning phase with unlabeled data, we reduce the learning weights of uncertain regions so that the model focuses more on learning predictions for certain regions. With this bidirectional two-stage learning strategy, we propose a new training method that can be directly applied to existing semi-supervised medical image segmentation methods. Extensive experiments show that our approach achieves new state-of-the-art (SOTA) performances for semi-supervised medical image segmentation, significantly improving different baseline performances.
We summarize our main contributions as follows:
@@
  • We propose a bidirectional uncertain region learning strategy. This strategy involves learning uncertain regions in labeled data as well as certain regions in unlabeled data so that the model can maximize the learning and correct the segmentation error regions.
@@
  • By focusing on high-uncertainty regions in labeled data, the model better utilizes label information to improve the prediction ability for difficult samples, while concentrating on certainty regions in unlabeled data helps minimize errors from incorrect pseudo-labels and stabilize training.
@@
  • The proposed method can be easily plug-and-play, which can be embedded into different methods to improve its performance. Extensive experiments validate the feasibility of the method, which shows significant improvement over previous SOTA methods on different datasets.
@@ 2.1 Semi-supervised Learning
Current semi-supervised learning methods [17] can generally be divided into two categories: consistency regularization [18, 19] and pseudo-labeling. [20, 21, 22].Pseudo-labeling refers to using the model’s predictions to generate labels for unlabeled samples, treating these predictions as approximate ”ground truth” to expand the training dataset and further train the model. Consistency regularization enforces the model’s outputs to be consistent for the inputs under different perturbations. In addition, some studies [23, 11] have combined the aforementioned methods and applied them to various tasks, showing superior performance.
@@ 2.2 Semi-Supervised Medical Image Segmentation
Semantic segmentation is the classification of every pixel in an image [24, 25]. It is a dense prediction task that requires substantial data and meticulous manual annotation in the training phase. Since semi-supervised methods can use unlabeled data to improve the performance of models, many semi-supervised methods have been applied to medical image segmentation in recent years. For example, SASSNet [26] introduces an edge distance function, calculating the signed distance from any voxel to its nearest natural boundary, and constructing a shape-aware semi-supervised semantic segmentation network. DTC [27] proposes a dual-task consistency framework, introducing dual-task consistency regularization between segmentation maps derived from level sets for labeled and unlabeled data and directly predicted segmentation maps. UA-MT [28], by incorporating uncertainty-aware consistency loss, guided by the estimated uncertainty, unreliable predictions are filtered out during the calculation of consistency loss, retaining only reliable predictions. However, removing unreliable predictions may result in the model ignoring features that are important but not easily detected, as well as meaning that they are underutilized, which may lead to loss of information.
@@ 2.3 Uncertainty Estimation
In semi-supervised learning, uncertainty estimates [29, 30, 31] can be used to assess the quality of model predictions. Methods for estimating uncertainty mainly include: (1) Using the information entropy of the predicted probability distribution to measure uncertainty, the higher the entropy value, the greater the uncertainty. (2) Utilizing the bias of multiple prediction results of the same inputs under different perturbations to measure uncertainty, the higher the bias, the higher the uncertainty of the model’s prediction of the input  [28]. (3) Calculate the variance of different prediction results for the same input to measure the uncertainty, the larger the variance, the higher the uncertainty of the model’s prediction for that input [32]. Since entropy is calculated based on the entire probability distribution, it is insensitive to fluctuations in individual predictions, which means that the entropy can reflect the overall uncertainty of the model more stably, rather than being overly affected by individual anomalous predictions. On the other hand, calculating the entropy value only requires the already obtained predictive probability distributions without additional model inference or complex calculations. Therefore, in our approach, we perform uncertainty estimation utilizing the information entropy of the predicted probability distribution.
3 Method
@@ 3.1 Preliminary
In semi-supervised segmentation, we assume that a training dataset $D$ contains $N$ labeled data and contains $M$ unlabeled data ($N\ll M$). The two are denoted as two subsets for convenience: $D=\mathcal{D}^{l}\cup\mathcal{D}^{u}$. Given a small labeled dataset $\mathcal{D}^{l}=\{(\mathbf{X}^{l}_{i},\mathbf{Y}^{l}_{i})\}_{i=1}^{N}$ and a large number of unlabeled images $\mathcal{D}^{u}=\{\mathbf{X}^{u}_{i}\}_{i=N+1}^{N+M}$. Where, $\mathbf{X}^{l}_{i}$ represents a labeled image, and $\mathbf{Y}^{l}_{i}$ represents its corresponding label. Similarly, $\mathbf{X}^{u}_{i}$ denotes an unlabeled image. Semi-supervised semantic segmentation aims at mining the unlabeled images with the help of limited labeling under effective supervision to obtain segmentation performance comparable to the results of the same type of segmentation with full supervision.
@@ 3.2 Overview
The overall pipeline of the proposed bidirectional uncertainty-aware region learning method is shown in Fig. 2, based on the teacher-student model [33] with a teacher network ${F}_{t}(\theta_{t})$ and a student network ${F}_{s}(\theta_{s})$. Due to the differences in the characteristics of uncertainty regions across different data, the identification and utilization of these uncertainty regions also vary. Therefore, we explore the scenarios of labeled and unlabeled data from two directions. For labeled images, we first use the teacher network for training. To enhance the model’s ability to learn potentially uncontrollable regions, we utilize the proposed URL. A mask is generated by calculating the entropy matrix of the model’s prediction, which is then used to divide the prediction into reliable and unreliable parts. Meanwhile, we supervise the same masking process for the ground-truth to ensure that the model learns mainly the uncertainty regions during training. For unlabeled data, we use the student network for training and use pseudo-labels generated by a teacher network trained from labeled data as supervision. To prevent the negative impact of potentially uncontrollable regions on the model, we utilize the proposed CRL with the same masking treatment for the model’s predictions and pseudo-labels to ensure that the model mainly learns predictions in deterministic regions during training, thus improving the stability and accuracy of the model.
@@ 3.3 Uncertainty-aware Region Learning
In semi-supervised medical image segmentation, there are inevitably potentially uncontrollable regions, and the presence of these regions increases the difficulty of model training. To allow the model to learn from potentially uncontrollable regions while maximizing the use of accurate annotation information, we force the model to learn potentially uncontrollable regions when training the model with labeled data. In this phase, we assign higher learning weights to these regions, allowing the model to better understand and correct prediction errors in these complex regions.
Specifically, we first train the teacher model using labeled data, for the uncertainty regions predicted by the segmentation model, the model has difficulty in making correct category judgments, to enable the model to better learn these challenging regions, we aim to enhance the learning intensity of uncertainty regions by leveraging the precise information of the label to improve the model’s predictive ability in these regions. We initially calculate the entropy matrix based on the model’s predictions. The entropy matrix measures the mapping of uncertainty corresponding to different category affiliations, and it reflects the overall uncertainty of the model for each predicted sample. If the entropy value is high, it indicates that the model is uncertain in its prediction for the region, and the predicted probabilities are relatively evenly distributed, with similar probabilities across categories, and cannot be categorized as belonging to a particular category. If the entropy value is very low, it implies that the model is highly certain in its prediction of the region, and the distribution of predicted probabilities is highly concentrated in a specific category. Utilizing the entropy matrix, we can distinguish between reliable and unreliable regions of model prediction.
@@
$logits^{l}_{i}=F_{t}(X^{l}_{i}),$ (1)
where $F_{t}$ is the teacher network that we need to pre-train. $logits^{l}_{i}$ are the logits outputs that correspond to $X^{l}_{i}$, Applying softmax to these logits yields the corresponding prediction probability scores:
$P^{l}_{i}=softmax(logits^{l}_{i}),$ (2)
where $P^{l}_{i}$ is the prediction of $X^{l}_{i}$. With the inference of prediction, uncertainty (i.e., entropy) maps corresponding to each class are computed according to the following equation:
@@
where $p$ is the output probability map of each class. To detect and learn the uncertain regions, we set a dynamic threshold and if the entropy value of a region exceeds this threshold, the region is considered as part of the highly uncertain region. Because labeled data contains accurate labels that can be directly utilized by the model for supervised learning, the model’s uncertain regions on labeled data can also be corrected by the label information. Therefore, a lower threshold can be set to include these regions more broadly, thus utilizing the label information more comprehensively in training. After finalizing the uncertainty region, we generate the corresponding image mask based on it $\mathbf{M}(\mathbf{M}\in\{0,1\}^{W\times H\times L})$. The formula is as follows:
@@
where $percentile(U,q)$ denotes the function that finds the $q$th percentile in the predicted probability distribution $U$ of the model each time. We then adjust the labeling of the labeled images utilizing the mask $M$, The formula is as follows:

### v2

@@ Abstract
In semi-supervised medical image segmentation, the poor quality of unlabeled data and the uncertainty in the model’s predictions lead to models that inevitably produce erroneous pseudo-labels. These errors accumulate throughout model training, thereby weakening the model’s performance. We found that these erroneous pseudo-labels are typically concentrated in high-uncertainty regions. Traditional methods improve performance by directly discarding pseudo-labels in these regions, which can also result in neglecting potentially valuable training data. To alleviate this problem, we propose a bidirectional uncertainty-aware region learning strategy to fully utilize the precise supervision provided by labeled data and stabilize the training of unlabeled data. Specifically, in the training labeled data, we focus on high-uncertainty regions, using precise label information to guide the model’s learning in potentially uncontrollable areas. Meanwhile, in the training of unlabeled data, we concentrate on low-uncertainty regions to reduce the interference of erroneous pseudo-labels on the model. Through this bidirectional learning strategy, the model’s overall performance has significantly improved. Extensive experiments show that our proposed method achieves significant performance improvement on different medical image segmentation tasks.
keywords:
Medical image segmentation, Semi-supervised learning, Uncertainty Learning.
journal: Information Fusion
@@ 1 Introduction
With the continuous development of deep learning, many medical image segmentation tasks [1, 2, 3] have achieved great success through fully supervised learning using large amounts of labeled data. However, it is challenging to obtain large-scale and accurately labeled medical datasets to train segmentation models due to the privacy, dispersion, and labeling difficulties of medical images [4]. Semi-supervised medical image segmentation [5, 6, 7, 8] effectively reduces the cost of labeled data and improves the performance and generalization ability of the model to a certain extent by utilizing a small amount of labeled data and a large amount of easily available unlabeled data to train the model.
Figure 1: Visualization of (a) pseudo-label error regions, (b) predicted high-entropy regions at the 99th percentile, (c) category probabilities for reliable predictions, and (d) category probabilities for unreliable predictions. The results show that most pseudo-label error regions strongly overlap with the model’s predicted high-entropy regions. Although unreliable predictions fluctuate across multiple categories, they can still effectively exclude certain categories (e.g., in (d) they exclude the background).
@@
In semi-supervised medical image segmentation, effectively leveraging unlabeled data is crucial. A common approach is assigning pseudo-labels to unlabeled data [9, 10], allowing the model to learn from these approximate “ground-truth”. However, generating high-quality pseudo-labels becomes particularly difficult due to the scarcity of labeled data. Low-quality pseudo-labels introduce greater uncertainty and contain numerous errors, which can mislead the model during training and ultimately affect its performance [11]. Therefore, effectively preventing the accumulation and propagation of errors in pseudo-labels has become a critical challenge in this area of research. Some researchers have proposed predictive filtering based on confidence scores to solve this problem. They argue that pseudo-tags generated by predictions with higher confidence scores are more reliable [12, 13, 14]. This approach aims to filter out low-confidence predictions and to retrain the model using only high-confidence pseudo-labels. However, the potential issue with relying solely on reliable predictions is that, while incorrect predictions are discarded, the model may never learn certain pixels throughout the training process, resulting in inadequate model training [15]. Particularly in medical images, each region may contain critical medical information [16]. If the model consistently to avoid these challenging regions during training, it will not only make it difficult to learn discriminative features for these regions, but will also cause the model’s performance on such regions to deteriorate, thus weakening the model’s ability to recognize critical structures. In addition, accurately defining whether an image is of high or low confidence is inherently challenging, further exacerbating this strategy’s instability.
To mitigate the issue of erroneous pseudo-labels more effectively, we analyze the model’s predicted entropy maps and the pseudo-label error maps obtained by comparing pseudo-labels with ground-truth labels, as shown in Fig. 1 (a) and (b), we find that incorrect pseudo-labels often occur in high-entropy regions, which largely coincide with the boundaries of the segmented targets. Therefore, we can effectively detect and correct pseudo-label errors by identifying and processing these high-entropy regions. Moreover, as shown in Fig. 1 (c) and (d), although the model’s predictions on the high-entropy regions are unreliable, it is very confident that these pixels do not belong to the background. Therefore, effectively utilizing this information can enhance the model’s learning. Based on the above analysis, we naturally identify two key challenges grounded in the characteristics of labeled and unlabeled data: (1) For labeled data, how can we leverage accurate ground-truth annotations to guide the correction of uncertain predictions. (2) For unlabeled data, in the absence of supervision, how can we avoid propagating errors while still extracting meaningful semantic cues from high-entropy regions?
To address the two challenges identified above, we propose a bidirectional region learning strategy tailored for semi-supervised segmentation, which consists of two complementary components: Uncertainty-aware Region Learning (URL) for labeled data and Certainty-aware Region Learning (CRL) for unlabeled data. (1) For labeled data, to correct predictions in high-uncertainty regions, we utilize the ground-truth to guide learning in these areas. Specifically, for each input image, we compute the prediction entropy map to quantify pixel-wise uncertainty. Based on the entropy values, predictions are categorized into high-entropy (uncertain) and low-entropy (certain) regions. During supervised training, we assign higher learning weights to high-entropy regions, encouraging the model to focus on difficult areas where errors are likely to occur. By leveraging precise supervision in these regions, the model gradually learns to identify and correct potential segmentation errors, thereby improving robustness and generalization. (2) For unlabeled data, in the absence of ground-truth supervision, we adopt a certainty-aware strategy to mitigate the impact of erroneous pseudo-labels in uncertain regions. Rather than discarding uncertain areas entirely (such as UA-MT calculates losses only in the deterministic region), we reduce their contribution to the loss by assigning lower learning weights during self-supervised learning. This soft suppression prevents noisy pseudo-labels from dominating training while still allowing the model to benefit from potentially informative features in a controlled manner. By integrating these two region-aware learning strategies, our approach not only guides the model to effectively learn the discriminative features of uncertain regions using the accurate Ground Truth and migrates them throughout the training process to enhance the model’s ability to recognize and model uncertain regions; at the same time, it also achieves robust learning of unlabeled data based on the reduction of noise sensitivity. Extensive experiments demonstrate that our approach achieves substantial performance gains across various baselines.
In summary, the contributions of this work consist of the following aspects:
  • 1.
@@
    We propose a new bi-directional uncertainty-aware learning strategy based on the difference between labeled and unlabeled data characteristics. This strategy involves learning uncertain regions in labeled data as well as certain regions in unlabeled data.
  • 2.
@@
    By focusing on the high-uncertainty regions in labeled data, the model’s ability to learn difficult samples can be improved; at the same time, focusing on the certainty regions in unlabeled data can help minimize the interference of erroneous pseudo-labels and stabilize the training.
  • 3.
@@
    The proposed method can be easily plug-and-play, which can be embedded into different methods to improve its performance. Extensive experiments validate the feasibility of the method.
@@ 2.1 Semi-supervised Learning
Current semi-supervised learning methods [17, 18] can generally be divided into two categories: consistency regularization [19, 20] and pseudo-labeling [21, 22, 23, 24]. Pseudo-labeling refers to using the model’s predictions to generate labels for unlabeled samples, treating these predictions as approximate ”ground truth” to expand the training dataset and further train the model. Consistency regularization enforces the model’s outputs to be consistent for the inputs under different perturbations. In addition, some studies [25, 11] have combined the methods above and applied them to various tasks, showing superior performance.
@@ 2.2 Semi-Supervised Medical Image Segmentation
Semantic segmentation is the classification of every pixel in an image [26, 27, 28]. It is a dense prediction task that requires substantial data and meticulous manual annotation in the training phase. To reduce the reliance on large-scale annotated data, many semi-supervised methods have been applied to medical image segmentation. For example, SASSNet [29] introduces an edge distance function, calculating the signed distance from any voxel to its nearest natural boundary, and constructing a shape-aware semi-supervised semantic segmentation network. DTC [30] proposes a dual-task consistency framework, introducing dual-task consistency regularization between segmentation maps derived from level sets for labeled and unlabeled data and directly predicted segmentation maps. UA-MT [31], by incorporating uncertainty-aware consistency loss, guided by the estimated uncertainty, unreliable predictions are filtered out during the calculation of consistency loss, retaining only reliable predictions. BCP [32] alleviates the experience mismatch between labeled and unlabeled data through bidirectional copy-paste operations. ABD [33] generates new samples based on adaptive bidirectional patch displacement guided by reliable prediction confidence, followed by training on these new samples to ensure the quality of consistent learning. However, due to the characteristics of medical images, directly removing unreliable regions may lead to the loss of critical anatomical structures, causing the model to overlook important yet difficult-to-detect features, thereby limiting its ability to learn from challenging samples.
@@ 2.3 Uncertainty Estimation
Uncertainty estimation [34, 35, 36, 37] has been widely explored in semi-supervised learning to calibrate model predictions or enhance the performance of prediction results. For example, Yu and Xia [31, 38] utilized Monte Carlo Dropout (MC-Dropout) [39] to estimate uncertainty, which was then employed to guide the model’s learning and promote the generation of more accurate pseudo-labels. However, MC-Dropout requires multiple forward passes, leading to significantly increased computational cost and time consumption. Zheng et al. [40] rectified noisy pseudo-labels by estimating uncertainty via the variance across multiple predictions of the same input. Luo et al. [41] refine the handling of uncertain regions by measuring the deviation between the mean prediction and the scaled prediction. Compared to MC-Dropout, these methods reduce computational overhead. Our method directly estimates uncertainty by calculating the entropy of the model’s predictive distribution, a process that does not require additional inference or complex computation and reflects the overall uncertainty stably, mitigating the disturbance of anomalous predictions.
3 Methodology
@@ 3.1 Preliminary
In semi-supervised segmentation, we assume that a training dataset $D$ contains $N$ labeled data and $M$ unlabeled data ($N\ll M$). For convenience, we represent them using sets: $D=\mathcal{D}^{l}\cup\mathcal{D}^{u}$. Given a small labeled dataset $\mathcal{D}^{l}=\{(\mathbf{X}^{l}_{i},\mathbf{Y}^{l}_{i})\}_{i=1}^{N}$ and a large number of unlabeled images $\mathcal{D}^{u}=\{\mathbf{X}^{u}_{i}\}_{i=N+1}^{N+M}$, where $\mathbf{X}^{l}_{i}$ represents a labeled image, and $\mathbf{Y}^{l}_{i}$ represents its corresponding label. Similarly, $\mathbf{X}^{u}_{i}$ denotes an unlabeled image. Semi-supervised semantic segmentation aims at mining the unlabeled images with the help of limited labeling under effective supervision to obtain segmentation performance comparable to the results of the same type of segmentation with full supervision.
@@ 3.2 Overview
The overall pipeline of the proposed bidirectional uncertainty-aware region learning method is shown in Fig. 2, based on the teacher-student model [42] with a teacher network ${F}_{t}(\theta_{t})$ and a student network ${F}_{s}(\theta_{s})$. Due to the differing characteristics of labeled and unlabeled data, our approach to identifying and utilizing uncertainty regions also differs. Specifically, for labeled images, we first train using the teacher network. To enhance the model’s ability to learn potentially uncontrollable regions, we use the proposed URL method, which generates a mask by calculating the entropy matrix of the model prediction and using it to divide both the model predictions and the corresponding ground truth into reliable and unreliable regions. We then assign higher learning weights to the unreliable regions, ensuring that the model focuses on uncertain areas during training. For unlabeled data, we use student networks for training and employ pseudo-labels generated by teacher networks trained on labeled data as supervision. To reduce the negative impact of potentially uncontrollable regions on the model, we apply the same mask treatment to the model predictions and pseudo-labels, and use the proposed CRL method to reduce the learning weights of unreliable regions, thus directing the model to mainly learn predictions from deterministic regions, improving the stability of training and the accuracy of the final segmentation.
@@ 3.3 Uncertainty-aware Region Learning
In semi-supervised medical image segmentation, regions that are difficult for the model to predict accurately inevitably exist due to factors such as blurry boundaries, similar tissue structures, or poor imaging quality. To allow the model to learn from potentially uncontrollable regions while maximizing the use of accurate annotation information, we assign higher learning weights to these regions during the training process with labeled data, guiding the model to focus on learning their feature distributions to enhance its understanding and prediction capability in complex regions.
Specifically, we first train the teacher model using labeled data and utilize the entropy matrix to distinguish between reliable and unreliable prediction regions in the labeled data. The entropy matrix reflects the model’s uncertainty in classifying each region. If the entropy value is high, it indicates that the prediction for the region is ambiguous, with similar probabilities across categories, making it difficult to determine the region’s category. Conversely, a low entropy value suggests that the model’s prediction for the region is highly confident, with the probability distribution concentrated on a specific category. Next, we leverage the precise information from the labels to enhance the learning intensity of the unreliable prediction regions, thereby improving the model’s prediction ability in these areas.
@@
$logits^{l}_{i}=F_{t}(\mathbf{X}^{l}_{i}),$ (1)
where $F_{t}$ is the teacher network that we need to pre-train. $logits^{l}_{i}$ are the logits outputs that correspond to $\mathbf{X}^{l}_{i}$. Applying softmax to these logits yields the corresponding prediction probability scores:
$\mathbf{P}^{l}_{i}=softmax(logits^{l}_{i}),$ (2)
where $\mathbf{P}^{l}_{i}$ is the prediction of $\mathbf{X}^{l}_{i}$. With the inference of prediction, uncertainty (i.e., entropy) maps corresponding to each class are computed according to the following equation:
@@
where $p$ is the output probability map of each class, that is, for the $i$-th labeled medical image, its value is $\mathbf{P}^{l}_{i}$. To detect and learn the uncertain regions, we set a dynamic threshold, and if the entropy value of a region exceeds this threshold, the region is considered part of the highly uncertain region. Because labeled data contains accurate labels that can be directly utilized by the model for supervised learning, the model’s uncertain regions on labeled data can also be corrected by the label information. Therefore, a lower threshold can be set to include these regions more broadly, thus utilizing the label information more comprehensively in training. After finalizing the uncertainty region, we generate the corresponding image mask based on it $\mathbf{M}(\mathbf{M}\in\{0,1\}^{W\times H\times L})$. The formula is as follows:
@@
where $percentile(U,q)$ denotes the function that finds the $q$-th percentile in the predicted probability distribution $U$ of the model each time. We then adjust the labeling of the labeled images utilizing the mask $\mathbf{M}$, The formula is as follows: